Slope of lines passing through points (1,2) and (3,4)
Answer by stanbon(75887)

Slope of line passing through points (1,2) and (3,4)
------
slope = (4-2)/(3-1) = 1
------
Form::
y = mx + b
-----
Solve for "b"::
2 = 1*1 + b
b = 1
---------
Equation:
y = x + 1
----------------
Cheers,
Stan H.
===============